calculate the molarity of each of the following solutions

1- 0.54 mole of BaCl2 in 100.0mL of solution

2- 0.230 mole of Na2CO3 in 200 mL

3- 0.325 mole of C6H12O6 in 250.0 mL

4- 1.384 mole of KNO3 in 250.0mL
Answer #1

Molarity has the equation:

M = mol of solute / Liter of solution

1)

M = mol/L = 0.54 / 100 ml = 0.54 / 0.1 L = 5.4 mol per liter

2)

M = mol/L = 0.23 / 200 ml = 0.23 / 0.2 L = 1.15 mol per liter

3)

M = mol/L = 0.325 / 250 ml = 0.325/ 0.25 L = 1.3 mol per liter

4)

M = mol/L = 1.384 / 250 ml = 1.384 / 0.25 L = 5.536 mol per liter

NOTE that mol per liter is equivalent to "Molar"